Vase I contains 2 white and 4 black balls and Vase II contains 4 white and 4 black
balls. If a ball is drawn out at random from one of the two vases, what is the
probability that it is a white ball?

Grade:12th pass

1 Answers

Arun
6 years ago
there will be two case for each vase
for vase I= 1/2*2/6= 1/6
for vase II=  1/2*4/6= 2/6
now the probability will be = 1/6+ 2/6= 3/6
Note- choosing a vase will also be a criteria hence multiplied by ½ in every case